OpenStudy (anonymous):

is the function f(x)=x^3+x even or odd or neither?

OpenStudy (anonymous):

odd

myininaya (myininaya):

look at the powers 3 and 1 3 and 1 are odd so the function is odd

myininaya (myininaya):

f(x)=x^2+1 is even f(x)=x^3+x^2 is neither since both exponents are not both even or both odd

OpenStudy (anonymous):

2^3 + 2 = 10 (-2)^3 + (-2) = -10 odd by definition if you put in (-x) you get -f(x) or if you put in -x you get the opposite of the answer than when you put in x

myininaya (myininaya):

yes you can use formal definition if you want it will be useful later if you plan on taking calculus

OpenStudy (anonymous):

thank yyouu

OpenStudy (anonymous):

does anybody know why a graph that is symmetric about the x axis cannot be the graph of a function

OpenStudy (anonymous):

use the vertical line rule if you draw a vertical line down the graph and you get 2 intersections, it's not a function by definition, the one x value of a function can only go to another y value different x values to can go to the same y value. but same x values cannot have different y values

OpenStudy (anonymous):

thankss how do i solve 3x^2+10x-8=0

OpenStudy (anonymous):

look at this site http://tutorial.math.lamar.edu/Classes/Alg/Factoring.aspx it teaches you have to factor look at example 3 part e